OpenStudy (anonymous):

PLEASE HELP !!!!!!!!!!!!!!!!!!!!!!!!!! A culture of bacteria contains 150 bacteria at time zero. Find the number of the bacteria. a) After 5 hours if their number doubles every hour. b) After 7 hours if their number triples every hour.

OpenStudy (anonymous):

doubles every hour first hour \(150\times 2\) second hour \(150\times 2^2\) third hour \(150\times 2^3\) etc

OpenStudy (anonymous):

Wouldn't we have to use Pert?

OpenStudy (anonymous):

no, not at all

OpenStudy (anonymous):

compute \[150\times 2^5\] is all

OpenStudy (anonymous):

Is there a specific formula for this?

OpenStudy (anonymous):

don't get married to formulas doubles every hour means... every hour it doubles in five hours it will double five times

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

yw